OpenStudy (anonymous):

Factor: x³ + 12x² + 48x + 64

OpenStudy (anonymous):

try f(-4) = -64 + 192 - 192 + 64 = 0 so (x+4) is a factor dividing f(x0 by x+4 gives (x^2 + 8x + 16) = (x+4)(x+4) f(x) = (x+4)^3

OpenStudy (anonymous):

get it marish? i used the factor theorem to find the first factor

OpenStudy (anonymous):

factor theorem: if f(x) is divisible by (x-a) then f(a) = 0 I had tried -1,-2 and -3 before i got to -4

OpenStudy (anonymous):

then i used long division to get the quadratic

OpenStudy (anonymous):

r u there?

OpenStudy (anonymous):

You could also realize that x^3+12x^2+48x+64 has x's in descending order (3,2,1,0) and that you can rewrite it as: x^3+3(4^1)x^2+3(4^2)x+1(4^3) This renders: (x+4)^3

OpenStudy (anonymous):

Yes, I understand. I was a little confused before, but thanks.